OpenStudy (anonymous):

PLZZ HELP Use the quadratic formula to find the zeros of the function y=17x^2-21

OpenStudy (goten77):

\[y=17x ^{2}-21\] well we dont have a b value so we ignore it and just plug in a,c and have \[(\pm \sqrt{-4\times 17 \times -21})/ 2 \times 17\] this becomes \[(\pm \sqrt{1428})/ 34\]

OpenStudy (goten77):

then we have \[\sqrt{1428} \approx 37.78\]

OpenStudy (goten77):

\[\pm37.78 / 34 \approx \pm 1.1\]

OpenStudy (anonymous):

but how is it -1.1 it should be just 1.1 cuz when u put -1.1 in it doesn't work

OpenStudy (goten77):

why would -1.1 not work? the out come becomes.... \[\pm37.78/34\]

OpenStudy (anonymous):

my calc messed up my bad